OpenStudy (anonymous):

What is the LCM of 35 and 25?

OpenStudy (anonymous):

\[35=5\times 7\\ 25=5\times 5\\ lcm=5\times 5\times 7\]

OpenStudy (anonymous):

It's easiest if you list the numbers out for both and then you can see it easily, for example for 25: 25, 50, 75...etc 35: 70, 105...etc and see which number will be in common

OpenStudy (anonymous):

The above method is very good as well! :)